Pls can you explain about relation between initial velocity and time t=0
Actually it is been said that initial velocity is the one which is at t=0
An example
If I watch a car moving in front of me and I start noting time , velocity etc
If I start noting velocity of car at 10.00
Am so is that 10 am is t=0 and if is yes
Then the further quest is if car was moving at 20 km /hr at 10.00 am
And at 10.01 am has 23km/hr velocity
And 10.05 am comes to rest
So what is it's initial and final velocity
Is initial velocity 20km/hr or 23km/hr
And what about final velocity is it 0
Pls clear this concept
1
Expert's answer
2019-05-06T10:21:11-0400

You are right: initial velocity is the one which is at t=0.

If you say that 0 am is t=0 then  initial velocity is 20km/hr:


"v_i=v(0)=20 \frac{km}{h}"

Final velocity will be the velocity at final time. If it is 10.01 am:


"v_f=23 \frac{km}{h}"